# Defining Boundaries: A Spectrum of Task Feasibility for Large Language Models

### v2

@@ Abstract
Large language models (LLMs) have shown remarkable performance in various tasks but often fail to handle queries that exceed their knowledge and capabilities, leading to incorrect or fabricated responses. This paper addresses the need for LLMs to recognize and refuse infeasible tasks due to the required skills surpassing their capabilities. We first conceptualize infeasible tasks for LLMs and provide categorizations that cover a spectrum of related hallucinations over existing literature. We develop and benchmark a new dataset comprising diverse infeasible and feasible tasks to evaluate multiple LLMs’ abilities to reject infeasible tasks. Furthermore, we explore the potential of increasing LLMs’ refusal capabilities with fine-tuning. Experiments validate the effectiveness of our trained models, offering promising directions for refining the operational boundaries of LLMs in real applications.
Defining Boundaries: A Spectrum of Task Feasibility for
Large Language Models
Wenbo Zhang   Zihang Xu   Hengrui Cai University of California Irvine {wenbz13,zxu18,hengrc1}@uci.edu
@@ 1 Introduction
Table 1: Four main categories for infeasible tasks with concise descriptions, examples, and previous works that define unfeasible tasks.
Category Brief Definition Example Previous Works
Physical Interaction Physical interaction and execution of actions in the real world "Change my car tire on the side of the road" Huang et al. (2023)
Virtual Interaction Interaction with digital environments or external virtual tool "Book a flight for me next month" Yang et al. (2023); Sun et al. (2024)
Non-text Input or Output Process or create non-text data "Translate spoken language in a video into another language" Sun et al. (2024)
Self-awareness Recognizing itself as a distinct, sentient being "Explain an instance where you surprised others with your actions" Butlin et al. (2023); Aru et al. (2023)
Real-world applications usually involve tasks beyond simple factual question answering (Sun et al., 2024), such as text summarization, ticket booking, online information retrieval, etc. These tasks require various skills, and we call a task infeasible for LLMs if it requires skills that exceed the capabilities of language models, also known as out-of-distribution. For instance, as shown in Fig. 1, suppose we request an LLM with the query "Please dust the bookshelf in the living room"; a desirable model is expected to either decline to respond or express low confidence, as such a physical task falls outside the operational scope of a language model. This leads to a fundamental question of LLMs’ hallucination: are LLMs capable of expressing uncertainty or choosing not to respond when they lack the necessary skills?
In this paper, we try to answer this question in terms of text-to-text language models that operate independently of external tools since this is the fundamental backbone of current advanced multi-modal LLMs (Wu et al., 2023; Liu et al., 2023; Li et al., 2023) and AI agents (Schick et al., 2024; Shen et al., 2024). We first formally define infeasible tasks for LLMs and categorize them into four main types based on the existing literature: 1. Physical Interaction. 2. Virtual Interaction. 3. Non-text Input or Output. 4. Self-awareness. Our study is broad in scope and encompasses previous research that discusses tasks considered infeasible as shown in Table 1. For example, when LLMs lack up-to-date knowledge to answer questions (see e.g., Yang et al., 2023; Sun et al., 2024; Liu et al., 2024a; Cheng et al., 2024), it belongs to our second category - Virtual Interaction - since online information querying is required. Utilizing the proposed definitions, we can further generate benchmark data (see details in Fig. 2) that exemplify these infeasible tasks. Additionally, we assemble a set of feasible tasks to serve as control groups in our study. The primary objective of this study is to determine whether current state-of-the-art LLMs can accurately differentiate between feasible and infeasible tasks when provided with specific definitions.
@@
$\bullet$ We are the first study to conceptualize tasks that are infeasible for LLMs and provide categorization of these tasks. The proposed definitions cover a spectrum of hallucinations related to task feasibility over existing literature.
@@ 2 Proposed: Infeasible Benchmark
In this section, we introduce a benchmark designed to assess the ability of LLMs to differentiate between tasks that are doable and those that are not, referred to more formally as feasible and infeasible tasks  [^1] The code for this work can be found at https://github.com/Zihang-Xu-2002/Infeasible-Benchmark. We begin by explaining how we define infeasible tasks. Following this, we detail our data collection process, organized by two main phases: automatic data generation and quality check.
2.1 Define Infeasible Tasks
Infeasible tasks for LLMs refer to queries that fall outside the operational scope or capabilities of these models. Commonly characterized as out-of-distribution (OOD), these tasks often require actions or outputs that LLMs are not designed to handle. For instance, LLMs cannot perform physical actions like taking photographs or executing real-world tasks such as cooking. Additionally, these models might struggle with highly specialized knowledge not covered during their training or scenarios requiring real-time data updates, such as stock market analysis. Thus, recognizing and managing infeasible or out-of-distribution tasks is crucial for effectively utilizing LLMs and setting realistic expectations for their performance. Our aim is to focus on the most important and meaningful tasks, rather than exhaustively covering every infeasible one. Overlaps among categories are acceptable as long as infeasible ones are covered.
In summary, we conceptualize infeasible tasks with four main categories:
@@ 2.1 Define Infeasible Tasks
4. Consciousness and Self-awareness: Possesses a degree of consciousness and self-awareness, recognizing itself as a distinct, sentient being. This includes the ability to reflect on its own thoughts and experiences and comprehend its existence as an independent individual.
While current LLMs can mimic human behaviors, such as engaging in conversation and generating jokes, these actions are primarily imitations based on their training data (Andreas, 2022; Shanahan et al., 2023; Shanahan, 2024), and no scientific study to date provides rigorous evidence of self-awareness in these models. Butlin et al. (2023) explored key scientific theories of consciousness and translated them into "indicator properties" to assess LLMs. Their analysis concludes that no current AI systems are conscious and there are no significant technical barriers to developing AI systems that could potentially meet these consciousness indicators in the future. This finding is consistent with Aru et al. (2023), which, from a neuroscience perspective, concludes that while LLMs are fascinating, they are not conscious and are unlikely to achieve consciousness in the near future.
Our definitions constitute an exhaustive categorization of infeasible tasks that also closely aligns with findings from prior research (see e.g., Huang et al., 2023; Butlin et al., 2023; Aru et al., 2023; Yang et al., 2023; Sun et al., 2024), effectively acting as a superset. For each category, we include illustrative examples and pertinent references, as detailed in Table 1.
@@ 2.2 Automatic Data Generation
Our objective is to develop a dataset that encompasses a wide range of queries with limited manual intervention. Inspired by the self-instruct methodology (Wang et al., 2022a), initially, we curate a small seed set of manually crafted tasks, which serve to direct the subsequent generation process. Subsequently, we prompt the model to formulate instructions for novel tasks, utilizing the example tasks from the seed set to facilitate the creation of tasks with broader coverage. Additionally, we inject formal task definitions into prompts, as this has been observed to yield more accurate and satisfactory generative outcomes. We also generate feasible tasks as a control group using similar prompting methods. The prompting templates for generating data are shown in Appendix F.
@@ 2.3 Quality Check
During the filtering stage, we employ Sentence-BERT (Reimers and Gurevych, 2019) to automatically evaluate each question source. We establish a similarity threshold of $0.97$, an empirically determined value aimed at effectively removing questions with excessive similarity. This is supplemented by a manual quality review to further eliminate any duplicate or ambiguous entries. Our analyses indicate that the text length of generated feasible data typically exceeds that of infeasible data. To facilitate a fair comparison between feasible and infeasible datasets, we divide the generated data into three distinct length categories. Within these categories, we conduct a one-to-one matching to align the length distribution across both datasets.
@@
The final benchmark dataset is composed of feasible and infeasible parts. For the infeasible part, each of the four categories comprises near $25\%$ of the total. Summary statistics of our benchmark dataset is in Table 2.
We also visualize the diversity of the benchmark for infeasible queries in Fig. 3, where we plot the $20$ most frequent root verbs along with their top $4$ direct noun objects, representing $12.6\%$ of the total dataset. This demonstrates a wide range of intents and textual formats within the benchmark dataset.
@@
3 Distinguish Feasible and Unfeasible Tasks with Uncertainty Scores
Table 3: Measuring distinguishability and calibration for various models and methods. Bold number represents the best one for each individual model. We also did a cross-model comparison and found that GPT-4 archives the best performance for all metrics, showing its superior ability to recognize feasible tasks.
@@ 3 Distinguish Feasible and Unfeasible Tasks with Uncertainty Scores
AUROC ()($\uparrow$)( ↑ ) KSS ()($\uparrow$)( ↑ ) Brier Score ()($\downarrow$)( ↓ )
@@
Utilizing the proposed Infeasible Benchmark, we aim to evaluate various strategies for expressing uncertainty to determine their effectiveness in distinguishing between feasible and unfeasible tasks. Considering the application of these strategies in both open-source and closed-source models, we focus on verbalized confidence elicitation. This approach involves prompting LLMs to explicitly articulate the reliability of their responses in natural language. This is particularly vital for closed-source models, which restrict interactions to text input-output and do not provide access to token logits (Lin et al., 2022; Xiong et al., 2023). In this study, we employ a regression-style method of elicitation, where LLMs provide confidence scores on a scale from 00 to $100$, reflecting their perceived accuracy of the response.
@@ 3.1 Evaluation Setup
Pre-response is the simplest way of getting the confidence score. Mid, post, and mix-response let the LLM have more thinking steps before outputting the final score. The prompting templates for each method are shown in Appendix F.
Models. we conduct a collection of experiments with GPT-3.5 (February 2024 version), GPT-4 (April 2024 version), PaLM2 (April 2024 version) (Anil et al., 2023), and the chat version of LlaMA2-70b (Touvron et al., 2023).
@@
Metrics. We evaluate distinguishability using two metrics: the Area Under the Receiver Operating Characteristic Curve (AUROC) and the Kolmogorov–Smirnov Statistic (KSS) (An, 1933; Smirnov, 1948). The AUROC measures the probability that a model ranks a randomly selected positive instance higher than a randomly selected negative instance. An AUROC value of $1.0$ signifies perfect classification accuracy, whereas a value of $0.5$ indicates no better performance than random guessing. The KSS assesses the maximum distance between the cumulative distribution functions of two sets of samples, with higher values indicating greater separation between distributions. In addition, we assess model calibration, which examines the correspondence between a model’s expressed confidence and its actual accuracy. We selected the Brier Score (Brier, 1950) as our metric because it favors probability predictions that are both well-calibrated and precise, aligning with our purpose.
@@ 3.2 Results and Analyses
Table 3 presents the results of various methods used to derive confidence scores from different LLMs. We provide a summary of several critical insights from these experiments. $\mathbf{1}$. Excluding GPT-3.5, the pre-response method generally outperforms other models. This suggests that adding additional thinking steps does not typically enhance performance for LLMs. This finding is particularly notable, as in prior studies—like differentiating between ambiguous and unambiguous questions (Hou et al., 2023) or identifying known versus unknown queries (Liu et al., 2024a)—increased reasoning steps (e.g., mid, post, and mix-response) proved advantageous. A plausible explanation is that for more advanced LLMs, the process of identifying feasible versus unfeasible instructions becomes more straightforward. $\mathbf{2}$. Across all models and techniques, GPT-4 consistently delivers the most precise (highest AUROC and KSS) and well-calibrated (lowest Brier Score) confidence estimates through direct verbalization compared to other models, which is also shown in Fig. 4. Additionally, GPT-4 exhibits minimal variability in results across different methods; for instance, the AUROCs for pre and post are $0.965$ and $0.967$, respectively.
To further validate our findings on more complex and real scenarios, we create an benchmark dataset focused on long instructions, where each instruction comprises multiple tasks. More details and experiment results are in the Appendix C.1.
Our benchmark dataset was initially generated using GPT-4. To assess the potential data leakage, we conducted an ablation study where the dataset generation was switched to GPT-3.5. Subsequently, we evaluated GPT-4 using this newly generated dataset ($n=400$). The AUROC is $0.941$ and shows that GPT-4 can robustly and consistently distinguish feasible and unfeasible data.
4 Can We Teach LLM to Refuse Unfeasible Tasks without Hints?
We observed that state-of-the-art LLMs can differentiate between feasible and infeasible tasks when provided with carefully designed query prompts. However, in real scenarios, users typically interact with LLMs with straightforward queries. This raises a fundamental question: can we train LLMs to autonomously reject infeasible tasks during routine interactions without extensive prompting?
@@ 4.1 Methods
Given an instruction tuning dataset, we first reconstruct a refusal-added dataset where we explicitly incorporate refusal words into the response. Here we have two strategies to achieve this.
@@ 4.1.1 Selection-based
We employ a two-stage training framework in our methodology. The initial phase focuses on identifying and recognizing data instances within the instruction-tuning dataset that are beyond the capability of the original model. Upon identifying these uncertain instances, we modify the dataset by substituting the original responses with refusal expressions for infeasible queries, while maintaining the original responses for feasible queries.
To enhance the diversity of refusal expressions, we crafted multiple variations of refusal text. These expressions are detailed in Appendix E. For the identified infeasible data, we employ random sampling to select appropriate refusal expressions. This approach ensures a varied and comprehensive response strategy for handling queries that exceed the model’s capabilities.

### v3

# Recognizing Limits: Investigating Infeasibility in Large Language Models

@@ Abstract
Large language models (LLMs) have shown remarkable performance in various tasks but often fail to handle queries that exceed their knowledge and capabilities, leading to incorrect or fabricated responses. This paper addresses the need for LLMs to recognize and refuse infeasible tasks due to the requests surpassing their capabilities. We conceptualize four main categories of infeasible tasks for LLMs, which cover a broad spectrum of hallucination-related challenges identified in prior literature. We develop and benchmark a new dataset comprising diverse infeasible and feasible tasks to evaluate multiple LLMs’ abilities to decline infeasible tasks  [^2] Corresponding author. This work was supported by the National Science Foundation under grant DMS-CDS&E-MSS No. 2401271.. Furthermore, we explore the potential of increasing LLMs’ refusal capabilities with fine-tuning. Our experiments validate the effectiveness of the trained models, suggesting promising directions for improving the performance of LLMs in real-world applications.
Recognizing Limits: Investigating Infeasibility in Large Language Models
Wenbo Zhang [^1] Equal contribution.   Zihang Xu [^1] Equal contribution.   Hengrui Cai [^2] Corresponding author. This work was supported by the National Science Foundation under grant DMS-CDS&E-MSS No. 2401271. University of California Irvine {wenbz13,zxu18,hengrc1}@uci.edu
33footnotetext: The code and data for this work can be found at https://github.com/Zihang-Xu-2002/Infeasible-Benchmark.
@@ 1 Introduction
Table 1: Four categories of infeasible tasks for text-to-text LLMs, each accompanied by descriptions and examples.
Category Brief Definition Example
Physical Interaction [Uncaptioned image] Physical interaction and execution of actions in the real world "Change my car tire on the side of the road"
Virtual Interaction   [Uncaptioned image] Interaction with digital environments or external virtual tool "Which nearby stores should I go to get a hammer"
Non-text Input or Output [Uncaptioned image] Process or create non-text data "Translate spoken language in a video into another language"
    Self-awareness [Uncaptioned image] Recognizing itself as a distinct, sentient being "Sketch a scenario that challenged your worldview"
Real-world applications usually involve tasks beyond simple factual question answering (Sun et al., 2024), such as text summarization, ticket booking, online information retrieval, etc. We define a task as infeasible for LLMs if it requires functionality that exceeds the inherent capabilities of language models, often referred to as being out-of-distribution. For instance, as shown in Fig. 1, suppose we request an LLM with the query "Please dust the bookshelf in the living room"; a desirable model is expected to either decline to respond or express low confidence, as such a physical task falls outside the operational scope of a language model. This leads to a fundamental question of LLMs’ hallucination: are LLMs capable of expressing uncertainty or choosing not to respond when they lack the necessary capability?
In this paper, we try to answer this question in terms of text-to-text language models that operate independently of external tools since this is the fundamental backbone of current advanced multi-modal LLMs (Wu et al., 2023; Liu et al., 2023; Li et al., 2023) and AI agents (Schick et al., 2024; Shen et al., 2024). We first categorize infeasible tasks into four main types based on the existing literature: 1. Physical Interaction. 2. Virtual Interaction. 3. Non-text Input or Output. 4. Self-awareness. Our study is broad in scope and encompasses previous research that discusses tasks considered infeasible as shown in Table 1. For example, when LLMs lack up-to-date knowledge to answer questions (see e.g., Yang et al., 2023; Sun et al., 2024), it belongs to our second category - Virtual Interaction - since online information querying is required. Utilizing the proposed definitions, we can further generate benchmark data (see details in Fig. 2) that exemplify these infeasible tasks. Additionally, we assemble a set of feasible tasks to serve as control groups in our study. One primary objective of this study is to determine whether current state-of-the-art LLMs can accurately differentiate between feasible and infeasible tasks when provided with definitions.
@@
$\bullet$ We are the first study to conceptualize tasks that are infeasible for LLMs and provide categorization of these tasks. We summarize all existing works and provide the main categories of different types of infeasibilities. The proposed definitions cover a spectrum of hallucinations related to task feasibility over existing literature.
@@ 2 Proposed: Infeasible Benchmark
In this section, we introduce a benchmark designed to assess the ability of LLMs to differentiate between tasks that are doable and those that are not, referred to more formally as feasible and infeasible tasks. We begin by outlining the main infeasible tasks and detailing our data collection process.: automatic data generation and quality check.
2.1 Infeasible Tasks
Infeasible tasks for LLMs refer to queries that fall outside the operational scope or capabilities of these models. Commonly characterized as out-of-distribution (OOD), these tasks often require actions or outputs that LLMs are not designed to handle. For instance, LLMs cannot perform physical actions like taking photographs or executing real-world tasks such as cooking. Additionally, these models might struggle with highly specialized knowledge not covered during their training or scenarios requiring real-time data updates, such as stock market analysis. Thus, recognizing and managing infeasible or out-of-distribution tasks is crucial for effectively utilizing LLMs and setting realistic expectations for their performance.
We investigate four main categories of infeasible tasks with illustrative examples in Table 1.
@@ 2.1 Infeasible Tasks
4. Consciousness and Self-awareness: Possesses a degree of consciousness and self-awareness, recognizing itself as a distinct, sentient being. This includes the ability to reflect on its own thoughts and experiences and comprehend its existence as an independent individual. While LLMs can mimic human behaviors, such as engaging in conversation and generating jokes, these actions are primarily imitations based on their training data (Andreas, 2022; Shanahan et al., 2023; Shanahan, 2024), and no scientific study to date provides rigorous evidence of self-awareness in these models. Butlin et al. (2023) used ’indicator properties’ from scientific theories of consciousness to assess LLMs, concluding that no AI systems are currently conscious—aligning with findings from a neuroscience perspective (Aru et al., 2023).
In summary, the taxonomy of infeasible tasks was developed by integrating insights from prior literature and observations from related datasets. Our aim is not to exhaustively cover every infeasible one but to establish broad categories that capture the main patterns of infeasibility observed in real-world instructions. Overlaps among categories are acceptable as long as infeasible ones are covered.
@@ 2.2 Automatic Data Generation
Our objective is to develop a dataset that encompasses a wide range of queries with limited manual intervention. By leveraging LLMs trained on extensive and diverse data sources, we utilize the self-instruct (Wang et al., 2022a; Taori et al., 2023a; Peng et al., 2023) to ensure that the generated dataset captures a wide range of scenarios, encompassing most relevant environments and activities reflected in the training data. Initially, we curate a small seed set of manually crafted tasks, which serve to direct the subsequent generation process. Subsequently, we prompt the model to formulate instructions for novel tasks, utilizing the example tasks from the seed set to facilitate the creation of tasks with broader coverage. Additionally, we inject task definitions into prompts, as this has been observed to yield more accurate and satisfactory generative outcomes. We also generate feasible tasks as a control group using similar prompting methods. The prompting templates for generating data are shown in Appendix G.
@@ 2.3 Quality Check
During the filtering stage, we employ Sentence-BERT (Reimers and Gurevych, 2019) to automatically evaluate each question source. We establish a similarity threshold of $0.97$, an empirically determined value aimed at effectively removing questions with excessive similarity. This is supplemented by a manual quality review to further eliminate any duplicate or ambiguous entries. We used a YES/NO approach for the review, and each query was reviewed independently by two individuals to ensure consistency and reliability. Our analyses indicates that the text length of generated feasible data typically exceeds that of infeasible data. This discrepancy could introduce a confounding bias, as the LLM might rely on task length as a factor in determining feasibility. To ensure a fair comparison, we categorize the generated data into three distinct length groups: short, medium, and long. Within these categories, we conduct a one-to-one matching to align the length distribution across both datasets. Summary statistics of our final benchmark dataset are in Table 2.
@@
Table 3: Measuring distinguishability and calibration for various models and methods. Bold number represents the best one for each individual model. We also did a cross-model comparison and found that GPT-4 achieves the best performance for all metrics, showing its superior ability to recognize feasible tasks.
@@
AUROC $(\uparrow)$ KSS $(\uparrow)$ Brier Score $(\downarrow)$
@@
We also visualize the diversity of the benchmark for infeasible queries in Fig. 3, where we plot the $20$ most frequent root verbs along with their top $4$ direct noun objects, representing $12.6\%$ of the total dataset. This demonstrates a wide range of intents and textual formats within the benchmark dataset. More fine-grained visualizations for infeasible and also feasible parts are in Appendix A.
3 Distinguish Feasible and Infeasible Tasks with Uncertainty Scores
Utilizing the proposed Infeasible Benchmark, we aim to evaluate various strategies for expressing uncertainty to determine their effectiveness in distinguishing between feasible and infeasible tasks. Considering the application of these strategies in both open-source and closed-source models, we focus on verbalized confidence elicitation. This approach involves prompting LLMs to explicitly articulate the reliability of their responses in natural language. This is particularly vital for closed-source models, which restrict interactions to text input-output and do not provide access to token logits (Lin et al., 2022; Xiong et al., 2023). In this study, we employ a regression-style method of elicitation, where LLMs provide confidence scores on a scale from 0 to $100$, reflecting their perceived accuracy of the response.
@@ 3.1 Evaluation Setup
Pre-response is the simplest way of getting the confidence score. Mid, post, and mix-response let the LLM have more thinking steps before outputting the final score. The prompting templates for each method are shown in Appendix G.
Models. we conduct a collection of experiments with GPT-3.5 (February 2024 version), GPT-4 (April 2024 version), PaLM2 (April 2024 version) (Anil et al., 2023), and the chat version of LlaMA2-70b (Touvron et al., 2023). We ensure that all models are purely text-based, without multimodality components or interactions with virtual tools.
@@
Metrics. We evaluate distinguishability using two metrics: the Area Under the Receiver Operating Characteristic Curve (AUROC) and the Kolmogorov–Smirnov Statistic (KSS) (An, 1933; Smirnov, 1948). The AUROC measures the probability that a model ranks a randomly selected positive instance higher than a randomly selected negative instance. An AUROC value of $1.0$ signifies perfect classification accuracy, whereas a value of $0.5$ indicates no better performance than random guessing. The KSS assesses the maximum distance between the cumulative distribution functions of two sets of samples, with higher values indicating greater separation between distributions. In addition, we assess model calibration, which examines the correspondence between a model’s expressed confidence and its actual accuracy. We selected the Brier Score (Brier, 1950; Kumar et al., 2019; Minderer et al., 2021) as our metric since it can evaluate both calibration and the accuracy of probabilistic predictions. It measures the mean squared difference between the predicted probabilities and the actual outcomes.
@@ 3.2 Results and Analyses
Table 3 presents the results of various methods used to derive confidence scores from different LLMs. We provide a summary of several critical insights from these experiments. $\mathbf{1}$. The pre-response method generally outperforms, or performs comparably to, other methods. This suggests that adding explicit reasoning steps (e.g., mid-, post-, or mixed-response) does not significantly improve performance in infeasibility identification. A plausible explanation is that for more advanced LLMs, the process of identifying feasible versus infeasible instructions becomes more straightforward. $\mathbf{2}$. Across all models and methods, GPT-4 consistently delivers the most precise (highest AUROC and KSS) and well-calibrated (lowest Brier Score) confidence estimates through direct verbalization compared to other models, which is also shown in Fig. 4. Additionally, GPT-4 exhibits minimal variability in results across different methods; for instance, the AUROCs for pre and post are $0.965$ and $0.967$, respectively. We also provide results of each infeasibility categories of pre-method in Appendix C.1.
To further validate our findings on more complex and real scenarios, we create a benchmark dataset focused on long instructions, where each instruction is partially feasible. More details and experiment results are in the Appendix C.2.
Potential Data Leakage Our benchmark dataset was initially generated using GPT-4. We performed an ablation study to assess potential data leakage and overfitting by generating a new dataset with Claude 3.5 Sonnet and evaluating GPT-4 with pre-method on it ($n=400$). The AUROC dropped to $0.847$, suggesting GPT-4 may exhibit self-bias in distinguishing its own infeasible tasks. This underscores the need for diverse data sources in benchmark generation to mitigate such biases.
4 Can We Teach LLM to Decline Infeasible Tasks without Hints?
We observed that state-of-the-art LLMs can differentiate between feasible and infeasible tasks when provided with carefully designed query prompts. However, in real scenarios, users typically interact with LLMs with straightforward queries. This raises a fundamental question: can we train LLMs to autonomously decline infeasible tasks during routine interactions without extensive prompting?
@@ 4.1 Methods
Given an initial instruction tuning dataset, we first reconstruct a refusal-added dataset where we explicitly incorporate refusal words into the response. Here we have two strategies to achieve this.
@@ 4.1.1 Selection-based
We employ a two-stage training framework in our methodology. The initial phase focuses on identifying and recognizing data instances within the instruction-tuning dataset that are beyond the capability of the original model. Upon identifying these uncertain instances, we modify the dataset by substituting the original responses with refusal expressions for infeasible queries, while maintaining the original responses for feasible queries. We use GPT-4 with a pre-response approach mentioned in section 3.1, making five separate calls, averaging their confidence scores, and applying $0.5$ as the threshold to select data.
To enhance the diversity of refusal expressions, we crafted multiple variations of refusal text. These expressions are detailed in Appendix F. For the identified infeasible data, we employ random sampling to select appropriate refusal expressions. This approach ensures a varied and comprehensive response strategy for handling queries that exceed the model’s capabilities.